OpenStudy (anonymous):

what is the answer to 3b-13+4b=7b+1

OpenStudy (anonymous):

3b-13+4b=7b+1 - always write out the problem first 7b-13=1 Combine like terms 7b=14 add 13 to both sides b=2 divide by 7

OpenStudy (anonymous):

\[3b-13+4b=7b+1\]combine the like terms:\[7b-13=7b+1\]add 13 to both sides:\[7b=7b+14\]subtract 7b from both sides:\[0\neq14\]thus there is no solution :)

OpenStudy (anonymous):

actually @julianstanley you left out a 7b, the equation is false :) @dmentedmonkey